OpenStudy (anonymous):

Find the least common denominator. Build up the fractions to equivalent fractions having the LCD as the denominator. 5/18 and 11/63

OpenStudy (asnaseer):

first what you need to do is find the smallest number that can be divided exactly by both 18 and 63. Can you do that?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

35/126 and 22/126

OpenStudy (asnaseer):

no. e.g. the smallest number that be be divided exactly by 8 and 6 is 24. because 24/8 = 3 and 24/6 = 4

OpenStudy (asnaseer):

abdul has basically done it for you, 126 is the smallest number that can be divided exactly by both 18 and 63. 18 * 7 = 126 63 * 2 = 126 then use the fact that any number multiplied by 1 is the same number, e.g.: x * 1 = x then use the fact that any fraction which has the same numerator and denominator is equal to 1, e.g.:\[\frac{a}{a}=1\]so you can then say:\[\frac{5}{18}=\frac{5}{18}*\frac{7}{7}=\frac{35}{126}\] use similar technique for the other fraction